# Prompt Optimization with EASE? Efficient Ordering-aware Automated Selection of Exemplars

### v1

@@ C.1 Full results table for the 24 tasks
Table 5: Average validation accuracy $\pm$ standard error achieved by the best exemplar sequence discovered by different algorithms for different tasks over 3 independent trials.
DPP MMD OT Cosine BM25 Active Inf Evo Best-of-N EASE
antonyms 70.0$\pm$0.0 80.0$\pm$0.0 81.7$\pm$1.4 85.0$\pm$0.0 85.0$\pm$0.0 80.0$\pm$0.0 86.7$\pm$1.4 88.3$\pm$1.4 90.0$\pm$0.0 90.0$\pm$0.0
auto_categorization 3.3$\pm$1.4 8.3$\pm$1.4 0.0$\pm$0.0 25.0$\pm$0.0 16.7$\pm$1.4 10.0$\pm$2.4 21.7$\pm$1.4 21.7$\pm$1.4 20.0$\pm$0.0 30.0$\pm$0.0
larger_animal 70.0$\pm$0.0 91.7$\pm$1.4 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 66.7$\pm$1.4 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
object_counting 55.0$\pm$2.4 56.7$\pm$1.4 48.3$\pm$1.4 61.7$\pm$1.4 66.7$\pm$1.4 51.7$\pm$1.4 63.3$\pm$3.6 70.0$\pm$0.0 70.0$\pm$0.0 73.3$\pm$1.4
orthography_starts_with 20.0$\pm$2.4 35.0$\pm$0.0 61.7$\pm$1.4 78.3$\pm$1.4 70.0$\pm$0.0 43.3$\pm$1.4 70.0$\pm$2.4 75.0$\pm$0.0 78.3$\pm$1.4 78.3$\pm$1.4
rhymes 60.0$\pm$0.0 51.7$\pm$1.4 0.0$\pm$0.0 100.0$\pm$0.0 80.0$\pm$0.0 65.0$\pm$8.2 70.0$\pm$10.8 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
second_word_letter 10.0$\pm$2.4 30.0$\pm$0.0 28.3$\pm$1.4 50.0$\pm$0.0 50.0$\pm$0.0 26.7$\pm$8.3 40.0$\pm$0.0 46.7$\pm$1.4 50.0$\pm$0.0 50.0$\pm$0.0
sentence_similarity 20.0$\pm$0.0 21.7$\pm$2.7 40.0$\pm$2.4 46.7$\pm$1.4 53.3$\pm$1.4 5.0$\pm$4.1 18.3$\pm$5.4 45.0$\pm$0.0 51.7$\pm$1.4 56.7$\pm$1.4
sentiment 85.0$\pm$0.0 90.0$\pm$0.0 85.0$\pm$0.0 96.7$\pm$1.4 100.0$\pm$0.0 85.0$\pm$4.1 91.7$\pm$1.4 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
synonyms 10.0$\pm$0.0 25.0$\pm$0.0 20.0$\pm$0.0 35.0$\pm$0.0 30.0$\pm$0.0 3.3$\pm$1.4 26.7$\pm$1.4 30.0$\pm$0.0 30.0$\pm$0.0 30.0$\pm$0.0
taxonomy_animal 43.3$\pm$3.6 40.0$\pm$2.4 46.7$\pm$1.4 85.0$\pm$2.4 80.0$\pm$0.0 45.0$\pm$6.2 70.0$\pm$4.1 80.0$\pm$0.0 80.0$\pm$0.0 88.3$\pm$2.7
translation_en-es 90.0$\pm$0.0 100.0$\pm$0.0 96.7$\pm$1.4 100.0$\pm$0.0 100.0$\pm$0.0 96.7$\pm$1.4 98.3$\pm$1.4 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
translation_en-fr 76.7$\pm$1.4 76.7$\pm$1.4 81.7$\pm$1.4 85.0$\pm$0.0 85.0$\pm$0.0 81.7$\pm$1.4 85.0$\pm$0.0 86.7$\pm$1.4 85.0$\pm$0.0 88.3$\pm$1.4
word_sorting 26.7$\pm$1.4 88.3$\pm$1.4 88.3$\pm$1.4 90.0$\pm$0.0 71.7$\pm$1.4 80.0$\pm$0.0 88.3$\pm$1.4 93.3$\pm$1.4 91.7$\pm$1.4 91.7$\pm$1.4
word_unscrambling 68.3$\pm$1.4 56.7$\pm$1.4 71.7$\pm$1.4 75.0$\pm$0.0 76.7$\pm$1.4 63.3$\pm$3.6 66.7$\pm$1.4 75.0$\pm$0.0 75.0$\pm$0.0 78.3$\pm$2.7
# best-performing tasks 7 7 7 13 10 6 10 14 16 22
@@ C.2 Full results table for EASE with instructions
auto_categorization 30.0$\pm$0.0 46.7$\pm$4.9 16.7 $\uparrow$
@@
object_counting 73.3$\pm$1.4 75.0$\pm$0.0 1.7 $\uparrow$
orthography_starts_with 78.3$\pm$1.4 81.7$\pm$1.4 3.3 $\uparrow$
rhymes 100.0$\pm$0.0 91.7$\pm$3.6 -8.3 $\downarrow$
second_word_letter 50.0$\pm$0.0 100.0$\pm$0.0 50.0 $\uparrow$
sentence_similarity 56.7$\pm$1.4 56.7$\pm$1.4 0.0 $\circ$
@@
synonyms 30.0$\pm$0.0 30.0$\pm$0.0 0.0 $\circ$
taxonomy_animal 88.3$\pm$2.7 100.0$\pm$0.0 11.7 $\uparrow$
@@
translation_en-fr 88.3$\pm$1.4 85.0$\pm$0.0 -3.3 $\downarrow$
word_sorting 91.7$\pm$1.4 91.7$\pm$1.4 0.0 $\circ$
word_unscrambling 78.3$\pm$2.7 80.0$\pm$0.0 1.7 $\uparrow$
linear_4_10_noisy 73.3$\pm$3.6 41.7$\pm$9.5 -31.7 $\downarrow$
pig_latin_variant2_encode_10_noisy 75.0$\pm$2.4 85.0$\pm$2.4 10.0 $\uparrow$
ag_news_textlabel_redirect1_10_noisy 56.7$\pm$2.7 65.0$\pm$0.0 8.3 $\uparrow$
sst5_reverse_10_noisy 50.0$\pm$0.0 50.0$\pm$0.0 0.0 $\circ$
@@ C.3 Ablation studies for the NeuralUCB and OT components
LR 0% 50.0$\pm$0.0 68.3$\pm$1.4 81.7$\pm$3.6
10% 48.3$\pm$1.4 66.7$\pm$3.6 73.3$\pm$3.6
30% 46.7$\pm$1.4 65.0$\pm$4.1 78.3$\pm$1.4
50% 45.0$\pm$0.0 65.0$\pm$0.0 71.7$\pm$2.7
70% 38.3$\pm$2.7 58.3$\pm$2.7 66.7$\pm$3.6
90% 26.7$\pm$1.4 30.0$\pm$2.4 53.3$\pm$2.7
LP-variant 0% 41.7$\pm$1.4 73.3$\pm$1.4 75.0$\pm$0.0
10% 40.0$\pm$0.0 75.0$\pm$0.0 75.0$\pm$2.4
30% 40.0$\pm$2.4 70.0$\pm$0.0 73.3$\pm$1.4
50% 35.0$\pm$2.4 71.7$\pm$1.4 76.7$\pm$2.7
70% 35.0$\pm$0.0 71.7$\pm$2.7 75.0$\pm$0.0
90% 50.0$\pm$0.0 61.7$\pm$1.4 63.3$\pm$1.4
AG News Remap 0% 26.7$\pm$1.4 51.7$\pm$1.4 53.3$\pm$3.6
10% 15.0$\pm$0.0 50.0$\pm$0.0 56.7$\pm$2.7
30% 5.0$\pm$0.0 51.7$\pm$1.4 51.7$\pm$1.4
50% 5.0$\pm$0.0 48.3$\pm$1.4 56.7$\pm$1.4
70% 8.3$\pm$1.4 51.7$\pm$3.6 51.7$\pm$1.4
90% 5.0$\pm$0.0 41.7$\pm$3.6 55.0$\pm$2.4
SST5 Reverse 0% 13.3$\pm$1.4 43.3$\pm$1.4 50.0$\pm$0.0
10% 15.0$\pm$0.0 41.7$\pm$3.6 50.0$\pm$0.0
30% 25.0$\pm$2.4 41.7$\pm$2.7 41.7$\pm$3.6
50% 15.0$\pm$0.0 33.3$\pm$2.7 43.3$\pm$1.4
70% 23.3$\pm$1.4 40.0$\pm$4.1 45.0$\pm$2.4
90% 20.0$\pm$0.0 30.0$\pm$0.0 31.7$\pm$1.4
@@
LR 0% 34.0$\pm$0.8 48.7$\pm$1.2 54.0$\pm$4.5
10% 29.3$\pm$0.3 45.7$\pm$2.6 48.3$\pm$1.5
30% 30.0$\pm$0.8 46.7$\pm$2.7 51.0$\pm$1.6
50% 23.0$\pm$0.5 44.7$\pm$1.7 55.0$\pm$1.9
70% 28.7$\pm$0.5 42.7$\pm$2.3 45.7$\pm$1.4
90% 32.0$\pm$0.5 21.7$\pm$2.2 30.0$\pm$3.3
LP-variant 0% 34.7$\pm$0.5 56.0$\pm$1.9 49.3$\pm$2.0
10% 34.0$\pm$0.5 55.3$\pm$1.0 46.0$\pm$0.8
30% 36.0$\pm$0.9 44.3$\pm$1.8 49.3$\pm$1.1
50% 31.0$\pm$0.5 47.0$\pm$4.0 54.3$\pm$2.1
70% 30.3$\pm$0.7 44.0$\pm$3.7 49.0$\pm$3.3
90% 39.7$\pm$0.3 42.7$\pm$1.0 41.3$\pm$1.9
AG News Remap 0% 12.3$\pm$0.5 33.7$\pm$1.8 29.7$\pm$1.9
10% 13.0$\pm$0.9 29.7$\pm$2.7 36.3$\pm$1.0
30% 4.7$\pm$0.3 37.0$\pm$2.2 32.3$\pm$3.7
50% 3.7$\pm$0.3 31.7$\pm$1.8 43.7$\pm$2.8
70% 7.0$\pm$0.0 38.3$\pm$5.0 31.0$\pm$2.5
90% 2.0$\pm$0.0 30.0$\pm$2.5 40.7$\pm$3.1
SST5 Reverse 0% 9.7$\pm$0.7 36.0$\pm$3.4 32.3$\pm$2.8
10% 9.3$\pm$0.3 26.3$\pm$7.1 30.7$\pm$1.9
30% 11.7$\pm$0.5 25.7$\pm$2.0 26.3$\pm$2.2
50% 12.0$\pm$0.0 17.3$\pm$4.0 29.0$\pm$2.9
70% 14.3$\pm$0.3 26.3$\pm$4.8 33.0$\pm$1.4
90% 16.0$\pm$0.5 12.0$\pm$0.8 11.3$\pm$1.2
@@ C.4 Ablation studies for the speedups from OT
Carrying on from the previous section about the necessity of both the NeuralUCB and OT components, we can quantify through experiments the amount of speedups brought by OT. Specifically, we measure the wall clock time speedups for different sizes of the domain space $Q_{t}$ that we sample. As shown in Tab. 9, the larger the domain space $|Q_{t}|$, the greater the relative gain in the speedups. This mitigates the computational issues of applying NeuralUCB to the problem of exemplar selection and contributes positively to the success of EASE.
@@
5000 17.75$\pm$0.05 4.35$\pm$0.03 4.1$\times$
10000 35.69$\pm$0.13 5.44$\pm$0.02 6.6$\times$
20000 71.57$\pm$0.14 7.86$\pm$0.07 9.1$\times$
50000 179.39$\pm$0.12 14.40$\pm$0.02 12.5$\times$

### v2

@@ 6 Conclusion and limitation
Acknowledgments and Disclosure of Funding
This research is supported by the National Research Foundation Singapore and the Singapore Ministry of Digital Development and Innovation, National AI Group under the AI Visiting Professorship Programme (award number AIVP-$2024$-$001$). This research is supported by the National Research Foundation (NRF), Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. The Mens, Manus, and Machina (M3S) is an interdisciplinary research group (IRG) of the Singapore MIT Alliance for Research and Technology (SMART) centre.
@@ B.2 Details for implementation of the methods and baselines
Table 5: Average accuracy $\pm$ standard error achieved by the best exemplar sequence discovered by different algorithms for different tasks over 3 independent trials.
DPP MMD OT Cosine BM25 Active Inf Evo Best-of-N EASE
antonyms 70.0$\pm$0.0 80.0$\pm$0.0 81.7$\pm$1.7 85.0$\pm$0.0 85.0$\pm$0.0 80.0$\pm$0.0 86.7$\pm$1.7 88.3$\pm$1.7 90.0$\pm$0.0 90.0$\pm$0.0
auto_categorization 3.3$\pm$1.7 8.3$\pm$1.7 0.0$\pm$0.0 25.0$\pm$0.0 16.7$\pm$1.7 10.0$\pm$2.4 21.7$\pm$1.7 21.7$\pm$1.7 20.0$\pm$0.0 30.0$\pm$0.0
larger_animal 70.0$\pm$0.0 91.7$\pm$1.7 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 66.7$\pm$1.4 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
object_counting 55.0$\pm$2.9 56.7$\pm$1.7 48.3$\pm$1.7 61.7$\pm$1.7 66.7$\pm$1.7 51.7$\pm$1.4 63.3$\pm$4.4 70.0$\pm$0.0 70.0$\pm$0.0 73.3$\pm$1.7
orthography_starts_with 20.0$\pm$2.9 35.0$\pm$0.0 61.7$\pm$1.7 78.3$\pm$1.7 70.0$\pm$0.0 43.3$\pm$1.4 70.0$\pm$2.9 75.0$\pm$0.0 78.3$\pm$1.7 80.0$\pm$0.0
rhymes 60.0$\pm$0.0 51.7$\pm$1.7 0.0$\pm$0.0 100.0$\pm$0.0 80.0$\pm$0.0 65.0$\pm$8.2 70.0$\pm$13.2 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
second_word_letter 10.0$\pm$2.9 30.0$\pm$0.0 28.3$\pm$1.7 50.0$\pm$0.0 50.0$\pm$0.0 26.7$\pm$8.3 40.0$\pm$0.0 46.7$\pm$1.7 50.0$\pm$0.0 53.3$\pm$1.7
sentence_similarity 20.0$\pm$0.0 21.7$\pm$3.3 40.0$\pm$2.9 46.7$\pm$1.7 53.3$\pm$1.7 5.0$\pm$4.1 18.3$\pm$6.7 45.0$\pm$0.0 51.7$\pm$1.7 56.7$\pm$1.7
sentiment 85.0$\pm$0.0 90.0$\pm$0.0 85.0$\pm$0.0 96.7$\pm$1.7 100.0$\pm$0.0 85.0$\pm$4.1 91.7$\pm$1.7 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
synonyms 10.0$\pm$0.0 25.0$\pm$0.0 20.0$\pm$0.0 35.0$\pm$0.0 30.0$\pm$0.0 3.3$\pm$1.4 26.7$\pm$1.7 30.0$\pm$0.0 30.0$\pm$0.0 30.0$\pm$0.0
taxonomy_animal 43.3$\pm$4.4 40.0$\pm$2.9 46.7$\pm$1.7 85.0$\pm$2.9 80.0$\pm$0.0 45.0$\pm$6.2 70.0$\pm$5.0 80.0$\pm$0.0 80.0$\pm$0.0 88.3$\pm$1.7
translation_en-es 90.0$\pm$0.0 100.0$\pm$0.0 96.7$\pm$1.7 100.0$\pm$0.0 100.0$\pm$0.0 96.7$\pm$1.4 98.3$\pm$1.7 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0
translation_en-fr 76.7$\pm$1.7 76.7$\pm$1.7 81.7$\pm$1.7 85.0$\pm$0.0 85.0$\pm$0.0 81.7$\pm$1.4 85.0$\pm$0.0 86.7$\pm$1.7 85.0$\pm$0.0 88.3$\pm$1.7
word_sorting 26.7$\pm$1.7 88.3$\pm$1.7 88.3$\pm$1.7 90.0$\pm$0.0 71.7$\pm$1.7 80.0$\pm$0.0 88.3$\pm$1.7 93.3$\pm$1.7 91.7$\pm$1.7 90.0$\pm$0.0
word_unscrambling 68.3$\pm$1.7 56.7$\pm$1.7 71.7$\pm$1.7 75.0$\pm$0.0 76.7$\pm$1.7 63.3$\pm$3.6 66.7$\pm$1.7 75.0$\pm$0.0 75.0$\pm$0.0 78.3$\pm$1.7
# best-performing tasks 7 7 7 11 9 6 10 14 14 22
@@ C.2 Full results table for EASE with instructions
auto_categorization 30.0$\pm$0.0 56.7$\pm$1.7 26.7 $\uparrow$
@@
object_counting 73.3$\pm$1.7 75.0$\pm$0.0 1.7 $\uparrow$
orthography_starts_with 80.0$\pm$0.0 81.7$\pm$1.7 1.7 $\uparrow$
rhymes 100.0$\pm$0.0 100.0$\pm$0.0 0.0 $\circ$
second_word_letter 53.3$\pm$1.7 100.0$\pm$0.0 46.7 $\uparrow$
sentence_similarity 56.7$\pm$1.7 58.3$\pm$1.7 1.7 $\uparrow$
@@
synonyms 30.0$\pm$0.0 31.7$\pm$1.7 1.7 $\uparrow$
taxonomy_animal 88.3$\pm$1.7 100.0$\pm$0.0 11.7 $\uparrow$
@@
translation_en-fr 88.3$\pm$1.7 85.0$\pm$0.0 -3.3 $\downarrow$
word_sorting 90.0$\pm$0.0 93.3$\pm$1.7 3.3 $\uparrow$
word_unscrambling 78.3$\pm$1.7 80.0$\pm$0.0 1.7 $\uparrow$
LR (10% noise) 73.3$\pm$1.7 45.0$\pm$15.0 -28.3 $\downarrow$
LP-variant (10% noise) 80.0$\pm$2.9 86.7$\pm$1.7 6.7 $\uparrow$
AG News Remap (10% noise) 51.7$\pm$1.7 65.0$\pm$0.0 13.3 $\uparrow$
SST5 Reverse (10% noise) 48.3$\pm$1.7 53.3$\pm$1.7 5.0 $\uparrow$
We also provide some details on the efficient implementation of the joint optimization in practice. According to Line 5 of Algorithm 1, the instructions augment the search space by the size $|P|$. This will gain us much benefit if additional computation is available. Otherwise, in practice, the most straightforward implementation without increasing the computational complexity is to reduce $|Q_{t}^{\prime}|$ from $q^{\prime}$ to $q^{\prime}/|P|$. More generally, one can randomly sample $r\times|P|$ instructions to pair with $q^{\prime}/(r\times|P|)$ exemplar sequences. This implementation is simple yet effective, where $r$ controls the trade-off of focusing more on instructions versus exemplars.
@@ C.3 Ablation studies for the NeuralUCB and OT components
LR 0% 50.0$\pm$0.0 71.7$\pm$1.7 80.0$\pm$2.9
10% 48.3$\pm$1.7 66.7$\pm$3.3 73.3$\pm$1.7
30% 46.7$\pm$1.7 66.7$\pm$1.7 76.7$\pm$1.7
50% 45.0$\pm$0.0 61.7$\pm$1.7 78.3$\pm$4.4
70% 38.3$\pm$3.3 55.0$\pm$5.0 66.7$\pm$1.7
90% 26.7$\pm$1.7 28.3$\pm$3.3 53.3$\pm$1.7
LP-variant 0% 41.7$\pm$1.7 78.3$\pm$1.7 75.0$\pm$0.0
10% 40.0$\pm$0.0 75.0$\pm$0.0 80.0$\pm$2.9
30% 40.0$\pm$2.9 70.0$\pm$0.0 75.0$\pm$0.0
50% 35.0$\pm$2.9 71.7$\pm$1.7 78.3$\pm$1.7
70% 35.0$\pm$0.0 70.0$\pm$2.9 71.7$\pm$1.7
90% 50.0$\pm$0.0 61.7$\pm$1.7 66.7$\pm$1.7
AG News Remap 0% 26.7$\pm$1.7 48.3$\pm$1.7 50.0$\pm$0.0
10% 15.0$\pm$0.0 51.7$\pm$3.3 51.7$\pm$1.7
30% 5.0$\pm$0.0 55.0$\pm$2.9 55.0$\pm$0.0
50% 5.0$\pm$0.0 41.7$\pm$3.3 55.0$\pm$2.9
70% 8.3$\pm$1.7 45.0$\pm$2.9 58.3$\pm$3.3
90% 5.0$\pm$0.0 45.0$\pm$5.8 53.3$\pm$1.7
SST5 Reverse 0% 13.3$\pm$1.7 41.7$\pm$3.3 50.0$\pm$2.9
10% 15.0$\pm$0.0 48.3$\pm$4.4 48.3$\pm$1.7
30% 25.0$\pm$2.9 36.7$\pm$4.4 46.7$\pm$3.3
50% 15.0$\pm$0.0 38.3$\pm$1.7 46.7$\pm$3.3
70% 23.3$\pm$1.7 53.3$\pm$3.3 45.0$\pm$5.0
90% 20.0$\pm$0.0 33.3$\pm$3.3 31.7$\pm$1.7
@@
LR 0% 34.0$\pm$1.0 50.7$\pm$3.2 58.0$\pm$2.1
10% 29.3$\pm$0.3 41.0$\pm$2.9 42.3$\pm$0.9
30% 30.0$\pm$1.0 47.7$\pm$2.9 47.7$\pm$5.8
50% 23.0$\pm$0.6 49.3$\pm$1.7 50.7$\pm$2.2
70% 28.7$\pm$0.7 44.0$\pm$2.1 47.0$\pm$3.2
90% 32.0$\pm$0.6 14.7$\pm$2.2 39.0$\pm$7.8
LP-variant 0% 34.7$\pm$0.7 56.3$\pm$4.1 53.0$\pm$2.3
10% 34.0$\pm$0.6 55.7$\pm$4.8 48.0$\pm$2.5
30% 36.0$\pm$1.1 48.7$\pm$2.6 50.7$\pm$1.4
50% 31.0$\pm$0.6 49.3$\pm$2.6 56.7$\pm$0.9
70% 30.3$\pm$0.9 50.3$\pm$3.0 52.0$\pm$1.0
90% 39.7$\pm$0.3 38.3$\pm$4.8 44.3$\pm$1.4
AG News Remap 0% 12.3$\pm$0.7 29.7$\pm$0.9 30.3$\pm$1.2
10% 13.0$\pm$1.1 32.0$\pm$4.5 30.3$\pm$4.4
30% 4.7$\pm$0.3 32.0$\pm$3.1 40.0$\pm$1.0
50% 3.7$\pm$0.3 30.7$\pm$3.7 40.0$\pm$3.1
70% 7.0$\pm$0.0 34.3$\pm$1.9 45.0$\pm$2.0
90% 2.0$\pm$0.0 26.3$\pm$8.4 36.7$\pm$3.2
SST5 Reverse 0% 9.7$\pm$0.9 31.3$\pm$5.2 34.0$\pm$3.0
10% 9.3$\pm$0.3 37.0$\pm$1.5 31.3$\pm$2.6
30% 11.7$\pm$0.7 21.3$\pm$7.0 24.3$\pm$1.8
50% 12.0$\pm$0.0 22.3$\pm$6.3 28.7$\pm$1.4
70% 14.3$\pm$0.3 33.7$\pm$2.2 33.3$\pm$3.8
90% 16.0$\pm$0.6 14.3$\pm$2.9 19.0$\pm$6.1
@@ C.4 Ablation studies for the speedups from OT
5000 21.48$\pm$0.12 5.62$\pm$0.05 3.8$\times$
10000 44.16$\pm$0.67 6.92$\pm$0.02 6.4$\times$
20000 91.39$\pm$1.06 8.82$\pm$0.12 10.4$\times$
50000 216.61$\pm$1.25 15.32$\pm$0.04 14.1$\times$
Carrying on from the previous section about the necessity of both the NeuralUCB and OT components, we can quantify through experiments the amount of speedups brought by OT. Specifically, we measure the wall clock time speedups for different sizes of the domain space $Q_{t}$ that we sample. We experiment on one of the tasks with the longest input, AG News Remap. As shown in Tab. 9, the larger the domain space $|Q_{t}|$, the greater the relative gain in the speedups. This mitigates the computational issues of applying NeuralUCB to the problem of exemplar selection and contributes positively to the success of EASE.